# ASGO: Adaptive Structured Gradient Optimization

### v2

@@ 6 Further Discussions on ASGO
preserving the matrix structure of the problem and naturally exploiting the structured properties because of the involvement of spectral norm in the steepest descent framework. This aligns with ASGO’s exploitation of structured properties. Moreover, if we ignore the momentum in both the gradient and the preconditioner, we can see that ASGO is equivalent to Muon. [^1] A proof of this equivalence can be found in Appendix G. Interestingly, this equivalence between ASGO and Muon in both theoretical basis and algorithm design is analogous to that between diagonal AdaGrad and SignSGD (Bernstein et al., 2018; Kunstner et al., 2023). In this sense, we may interpret ASGO and Muon as AdaGrad and SignSGD for structured gradients, respectively. Note that Shampoo also admits such a relation with Muon in algorithmic design. However, as discussed in Section 4, Shampoo has a worse convergence rate and thus fails to benefit from the structured properties as does ASGO. To some extent, this means that ASGO may be a more appropriate approach than Shampoo as such an analog to diagonal AdaGrad.
@@
However, in smooth settings, nonconvex convergence analysis of Muon has been presented in Li and Hong (2025). They establish the convergence results under a more general setting, involving gradient noise and momentum following the analysis in Cutkosky and Mehta (2020). However, if we only look at the deterministic case, their result is worse than (3) because of the explicit dependence on the dimension $n$. The key here is that the standard smoothness condition with respect to the Frobenius norm is not a good fit for analyzing structured gradient algorithms like Muon or ASGO. Using Assumption 2, we obtain better convergence results for Muon in (3).
@@
We also provide Algorithm 2, a practical implementation of ASGO inspired by Adam (Kingma and Ba, 2014) and distributed Shampoo (Shi et al., 2023), together with a memory efficient diagonal variant of ASGO, named DASGO in Algorithm 3. DASGO is basically a light-weight optimizer with the preconditioner $\Lambda_{t}$ of ASGO being diagonalized, which is efficient in both memory and computations. Please refer to Appendix B for details.
@@ 7 Empirical Results
We empirically evaluate the effectiveness of ASGO (Algorithm 2) and DASGO (Algorithm 3) on pretraining and finetuning tasks for Large Language (LL) models. We compare the methods against established optimizers, including AdamW (Kingma and Ba, 2014; Loshchilov and Hutter, 2017), Shampoo (Gupta et al., 2018), and Muon (Jordan et al., 2024). Several important implementational details should be noted for fair comparison. First, since Muon is designed to operate exclusively on matrix parameters, we follow Jordan et al. (2024) and apply AdamW update rules to all 1D parameters within the Muon optimizer to ensure that it can handle the complete model. All experiments were conducted using NVIDIA V100s SMX2 GPUs. Specifically, for the larger-scale pretraining of GPT2, we utilized a configuration of four V100 GPUs, while other experiments were performed on a single V100 GPU.
7.1 Pretraining NanoGPT
We first conducted experiments using the NanoGPT architecture on the Shakespeare character dataset, following the configuration in (Karpathy, 2022). To ensure a fair comparison, we maintained consistent hyperparameter (HP) settings varying only those parameters that significantly impacted optimizer performance, i.e., learning rate, 1st and 2nd order moment coefficients ($\beta_{1}$ and $\beta_{2}$) and so on)  [^2] A full description of how these “best” HPs were selected and their values are presented in Appendix C.. For consistency across all optimization methods, we employ the OneCycleLR learning rate schedule, which has been shown to provide stable convergence properties in deep learning tasks. Figure 1 presents a comparison of the pretraining performance of ASGO and DASGO with that of Shampoo, Muon and AdamW on the NanoGPT model. See Appendix C for a plot of the training loss.
Figure 1: Validation Loss for NanoGPT
Examining both test and training loss curves, ASGO consistently outperforms Shampoo despite requiring only half the memory consumption and computational effort, which highlights ASGO’s practical advantages for training language models. Furthermore, ASGO and Muon achieve the lowest final training and test losses, outperforming all other methods. This consistent performance between ASGO and Muon aligns well with our discussions in Section 6. Moreover, The lightweight DASGO optimizer achieves competitive results against AdamW in both training and test loss metrics. DASGO demonstrates particularly strong performance during the initial training phase (first 200 training steps, which is about the first two epochs), but eventually exhibits a performance gap with ASGO. We will discuss this performance gap in Section 7.2.
@@ 7.2 Pretraining GPT2
To further evaluate the efficacy of ASGO and DASGO, we extended our investigation to larger-scale pretraining tasks. We adopted the configuration of the GPT2 model as described by (Karpathy, 2022). In this setting, the GPT2 model comprises 12 Transformer blocks, each with a hidden dimension of 768. Consequently, there are 12 projection layers with dimensions 768×2304. In each of these layers alone, Shampoo necessitates storing two matrices of size 768×768 and two matrices of size 2304×2304 for its state. In contrast, ASGO requires storing only two 768×768 matrices, corresponding to the preconditioning and inverse preconditioning factors. This substantial difference in memory requirements rendered Shampoo impractical for our large-scale GPT2 pretraining experiments, preventing us from obtaining meaningful results within our computational constraints. Therefore, we do not report Shampoo’s performance in this setting. To ensure a fair comparison, we carefully tuned the learning rates and $\beta_{2}$ values (where applicable) for all optimizers (tuning details are provided in Section C). Furthermore, we employed a learning rate schedule consisting of 200 linear warm-up steps followed by a cosine decay, which is a common practice for training LLM models. Figure 2a and Table 1 present a comparison of the training and validation loss achieved by ASGO and DASGO against Muon and AdamW on the GPT2 model.
Table 1: Pretraining GPT2 Validation Loss
AdamW Muon ASGO DASGO
3.96 3.82 3.87 5.23
The training loss curves reveal that ASGO achieves a lower final training loss compared to AdamW in the pretraining task, suggesting superior optimization. This observation is further corroborated by the validation loss (Table 1), where ASGO’s performance (3.87) is better than AdamW’s (3.96). Interestingly, ASGO’s performance appears more similar to Muon (3.82) in terms of final loss, although ASGO exhibits a slightly higher validation loss than Muon.
(a) Train Loss
(b) Learning Rate Sensitivity
Figure 2: Pretraining Performance and Learning Rate Sensitivity on GPT2
@@
ASGO demonstrates potential advantages over Muon. As depicted in Figure 2a, ASGO shows a faster initial decrease in training loss. Furthermore, our learning rate sensitivity analysis (Figure 2b), with tested candidates $[10^{-3},5\times 10^{-3},10^{-2},3\times 10^{-2},5\times 1e^{-2},7\times 10^{-2},10^{-1},5\times 10^{-1}]$ using optimal settings for Muon and ASGO suggests that ASGO might be more robust to learning rate selection. Specifically, at higher learning rates (e.g., $10^{-1}$, $5\times 10^{-1}$), ASGO maintains relatively stable training while Muon exhibits instability. These observations – faster initial convergence and a potentially wider stable learning rate regime for ASGO – are consistent with characteristics often associated with adaptive optimization methods, as discussed in Section 6. However, DASGO (5.23) underperformed compared to AdamW, Muon, and ASGO in the GPT2 pretraining task. This performance gap was also observed in the pretraining of the smaller NanoGPT model. A primary reason for the difference between DASGO and ASGO likely stems from its diagonal preconditioning. By only retaining and using only the diagonal elements of the $GG^{T}$, DASGO essentially disregards the inter-dependencies between different parameter gradients from neurons within a layer, moving away from a true matrix-based adaptive approach towards a per-parameter scaling akin to vector-based methods. This highlights the value of preserving non-diagonal elements in the preconditioner matrices, particularly for capturing parameter interactions in attention-based architectures. Nevertheless, DASGO’s significantly reduced memory footprint makes it a compelling option for resource-limited settings where computational efficiency is paramount.
@@ 7.3 Finetuning GPT2-Large on WikiText-2
To account for statistical variability, we conducted five experimental runs with different random seeds under the same hyperparameter settings. Table 2 presents the average perplexity results after fine-tuning for 2 epochs. For both objectives, we employed a cosine decay learning rate scheduler. Hyperparameter search details and 95% confidence intervals are provided in Table 6.
@@
Under the CLM objective, ASGO (13.88 perplexity) and DASGO (13.84 perplexity) achieved lower perplexity than both Muon (13.91 perplexity) and AdamW (14.01 perplexity). This suggests that both ASGO and its memory-efficient variant, DASGO, can be effective for traditional LL modeling fine-tuning. When the model was fine-tuned using the FIM objective, we observed a similar trend, with ASGO (15.66 perplexity) and DASGO (15.45 perplexity) outperforming both Muon (15.93 perplexity) and AdamW (17.46 perplexity). Across both fine-tuning scenarios, ASGO and DASGO demonstrated competitive or superior performance compared to these baseline optimizers.

### v3

@@ 6 Further Discussions on ASGO
preserving the matrix structure of the problem and naturally exploiting the structural properties because of the involvement of spectral norm in the steepest descent framework. This aligns with ASGO’s exploitation of the structured properties of the gradient. Moreover, if we ignore the momentum in both the gradient and the preconditioner, we can see that ASGO is equivalent to Muon. [^1] A proof of this equivalence can be found in Appendix G. Interestingly, this equivalence between ASGO and Muon in both theoretical basis and algorithm design is analogous to that between diagonal AdaGrad and SignSGD (Bernstein et al., 2018; Kunstner et al., 2023). In this sense, we may interpret ASGO and Muon as AdaGrad and SignSGD for structured gradients, respectively. Note that Shampoo also admits such a relation with Muon in algorithmic design. However, as discussed in Section 4, Shampoo has a worse convergence rate and thus fails to benefit from the structured properties as does ASGO. To some extent, this means that ASGO may be a more appropriate approach than Shampoo as such an analog to diagonal AdaGrad.
@@
In smooth settings, nonconvex convergence analysis of Muon has been presented in Li and Hong (2025). These convergence results are established under a more general setting, involving gradient noise and momentum following the analysis in Cutkosky and Mehta (2020). However, if we only look at the deterministic case, their result is worse than (3) because of the explicit dependence on the dimension $n$. The key here is that the standard smoothness condition with respect to the Frobenius norm is not a good fit for analyzing structured gradient algorithms like Muon or ASGO. Using Assumption 2, we obtain better convergence results for Muon in (3).
@@
We also provide Algorithm 2, a practical implementation of ASGO, together with a memory-efficient diagonal variant of ASGO, named DASGO in Algorithm 4. DASGO is basically a light-weight optimizer with the preconditioner $\Lambda_{t}$ of ASGO being diagonalized, which is efficient in both memory and computations. To make ASGO effective and efficient in practice, a critical aspect is the efficient computation of the preconditioner’s inverse square root, $V_{t}^{-1/2}$. Standard methods such as Singular Value Decomposition (SVD) are prohibitively expensive for large-scale applications. To ensure our optimizer is practical, we adopt the Newton-Schultz (NS) iterative method as summarized in Algorithm 3, which substantially reduces the wall-clock time per step. The applicability of this approach stems from the deep connection between the matrix inverse square root and the matrix sign function. As established in Thm 5.2 (Higham, 2008b), computing $A^{-1/2}$ is equivalent to extracting a sub-block from the matrix sign function of an augmented matrix:
$\operatorname{msign}\left(\left[\begin{array}[]{cc}0&A\\ I&0\end{array}\right]\right)=\left[\begin{array}[]{cc}0&A^{1/2}\\ A^{-1/2}&0\end{array}\right].$
This equivalence implies that the computational complexity of our inverse square root step is on the same order as the matrix operation in the Muon optimizer. It is worth noting, however, that the NS iteration in ASGO is applied to this $2\times 2$ block matrix, leading to a slightly heavier per-iteration cost compared to Muon. Finally, we highlight that our framework is modular. Recent advances (Amsel et al., 2025; Grishina et al., 2025) compute the matrix sign function more efficiently than classical Newton-Schultz, for instance, by optimizing the iteration’s hyperparameters. This modularity allows such methods to be readily incorporated. In our empirical implementation, we validated this by experimenting with both the classical Newton-Schultz and PolarExpress (Amsel et al., 2025) methods. Please refer to Appendix B for more details of other modifications in our practical implementation.
@@ 7 Empirical Results
We empirically evaluated the effectiveness of ASGO (Algorithm 2) and DASGO (Algorithm 4) on pretraining and finetuning tasks for Large Language models (LLMs). We compared our methods against the established optimizers, AdamW (Kingma and Ba, 2014; Loshchilov and Hutter, 2017), Shampoo (Gupta et al., 2018), and Muon (Jordan et al., 2024). Several important implementational details should be noted for fair comparison. First, since Muon is designed to operate exclusively on matrix parameters, we followed Jordan et al. (2024) and applied AdamW update rules to all 1D parameters within the Muon optimizer to ensure that it can handle the complete model. All experiments were conducted using NVIDIA V100s SMX2 and NVIDIA GH200 GPUs. Specifically, for the larger-scale pretraining of GPT2, we utilized a configuration of four GH200 GPUs, while other experiments were performed on a single V100 GPU.
@@ 7.1 Pretraining GPT2
[…]
Table 1: Pretraining GPT2 Train loss and Validation Loss
Optimizer Training Loss Validation Loss
AdamW 3.442 3.455
Muon 3.332 3.342
Shampoo 3.405 3.417
DASGO (Ours) 3.497 3.509
ASGO-NewtonSchultz (Ours) 3.335 3.347
ASGO-PolarExpress (Ours) 3.332 3.342
We now present the performance evaluation of ASGO against several baselines on the GPT-2 pretraining task. The final training and validation losses after 2400 steps are summarized in Table 1, and the full training loss dynamics are depicted in Figure 1.
Figure 1: Pretraining GPT-2 Train Loss.
Although the results obtained using ASGO-PE and ASGO-NS are similar, ASGO-PE performs slightly better than ASGO-NS. These results clearly demonstrate that ASGO achieves state-of-the-art performance, matching the strongest baseline while significantly outperforming AdamW and Shampoo. Focusing on the final validation loss in Table 1, ASGO-PolarExpress achieves a validation loss of 3.342. This result is on par with the Muon optimizer (3.342). More importantly, ASGO’s performance is substantially better than both the standard AdamW (3.455) and Shampoo (3.417).
The training curves in Figure 1 provide deeper insights into the training dynamics. ASGO and Muon exhibit nearly identical training trajectories; both converge faster than AdamW. In contrast, while Shampoo displays loss spikes during the training process. This phenomenon suggests potential numerical instability even with preconditioner update frequency 1, a behavior not observed in ASGO.
These findings strongly support the efficacy of ASGO. It not only matches the final performance and rapid convergence of the highly-tuned Muon optimizer but also remedies the training instabilities observed in Shampoo.
@@
However, DASGO (3.509) underperformed AdamW, Muon, Shampoo and ASGO in the GPT2 pretraining task. This performance gap was also observed in the pretraining of the smaller NanoGPT model. A primary reason for the difference between DASGO and ASGO likely stems from its diagonal preconditioning. By only retaining and using only the diagonal elements of the $GG^{T}$, DASGO essentially disregards the inter-dependencies between different parameter gradients from neurons within a layer, moving away from a true matrix-based adaptive approach towards a per-parameter scaling akin to vector-based methods. This highlights the value of preserving non-diagonal elements in the preconditioner matrices, particularly for capturing parameter interactions in attention-based architectures. Nevertheless, DASGO’s significantly reduced memory footprint makes it a compelling option for resource-limited settings where computational efficiency is paramount.
7.1.1 Wall time comparison for pretraining GPT2
We now compare the Wall-Clock Times (WCTs) per training step for AdamW, Muon and ASGO, using the exact experimental setup from the GPT-2 pretraining task (Section 7.1).
Table 2: Wall-clock time per training step, difference compared with AdamW update.
Optimizer Single Train Step WCT Time Difference
AdamW 5.451 sec 0 sec
Muon 5.519 sec 0.068 sec
ASGO (Ours) 5.947 sec 0.468 sec
For this analysis, we used 10 iterations of the NS algorithm for ASGO instead of 5 iterations used for Muon’s update, which are the settings required to achieve the optimal convergence performance reported previously. The measured WCT for a single training step (including forward, backward, gradient accumulation passes, and optimizer steps) is shown in Table 2. Although computing the inverse square root of $V_{t}$, in ASGO’s optimizer step is more expensive than in AdamW, the forward and backward passes remain the dominant computational bottleneck. As shown in Table 2, a full training step with ASGO costs only 0.468 seconds more than the AdamW update. This overhead (8.6% increase over AdamW) is acceptable, and consequently, the total end-to-end training time is only marginally affected.
These tables show that although AdamW is computationally cheapest at the optimizer level, the total training times for all three methods are very similar. This demonstrates that the total training time may not be a main concern for ASGO. Note that ASGO achieves the smallest training loss while AdamW achieves the largest loss.
7.1.2 Ablation Study on the preconditioning side:
In Section 5, we theoretically motivated our choice of applying the single-sided preconditioner to the right side of the gradient (or its momentum), based on block-wise smoothness convergence results. To empirically validate this crucial design choice, we conducted an ablation study comparing the performance of ASGO with left-sided versus right-sided preconditioning. The experimental setup is identical to that used for the GPT-2 pretraining in Section 7.1. As presented in Table 3, the results demonstrate that right-sided preconditioning consistently and significantly outperforms its left-sided counterpart across all tested learning rates ({5e4,1e3,5e3,1e25e-4,1e-3,5e-3,1e-2}).
This finding empirically confirms our theoretical intuition and highlights the clearly asymmetric impact of the preconditioning side. This observed asymmetry also suggests a potential limitation of Shampoo-like double-sided preconditioners. Beyond their higher memory and computational costs, their symmetric approximation (derived from a Kronecker product) may fail to capture the true underlying structured curvature. This could lead to misplacing the preconditioning effect on the less critical side, resulting in less effective preconditioning. We hypothesize that this structural mis-approximation could be a contributing factor to the sub-optimal performance and training instabilities (e.g., the loss spike in Figure 1) observed in Shampoo.
Table 3: Performance comparison between precondition side
Learning Rate Validation Loss of left side Validation Loss of right side
5e-4 3.582 3.569
1e-3 3.468 3.447
5e-3 3.399 3.362
1e-2 3.412 3.386
@@ 7.2 Finetuning GPT2-Large on WikiText-2
To account for statistical variability, we conducted five experimental runs with different random seeds under the same hyperparameter settings. Table 4 presents the average perplexity results after fine-tuning for 2 epochs. For both objectives, we employed a cosine decay learning rate scheduler. Hyperparameter search details and 95% confidence intervals are provided in Table 8 in Appendix C3.
@@
Under the CLM objective, ASGO (13.88 perplexity) and DASGO (13.84 perplexity) achieved lower perplexity than both Muon (13.91 perplexity) and AdamW (14.01 perplexity). This suggests that both ASGO and its memory-efficient variant, DASGO, can be effective for traditional LLM fine-tuning. When the model was fine-tuned using the FIM objective, we observed a similar trend, with ASGO (15.66 perplexity) and DASGO (15.45 perplexity) outperforming both Muon (15.93 perplexity) and AdamW (17.46 perplexity). Across both fine-tuning scenarios, ASGO and DASGO demonstrated competitive or superior performance compared to these baseline optimizers.